OpenStudy (anonymous):

Find the pattern and the fourth term: 1/1, 1/4, 1/9 HELP!!!!!!

OpenStudy (freckles):

hint 1^2=1 2^2=4 3^2=9

OpenStudy (anonymous):

the fourth term is 1/16

OpenStudy (campbell_st):

@isabelandscrappy thanks for providing the answer, but one of the objectives of the site is to help people improve their understanding... answers don't actually do that.